Sin1sin2sin3_______sin179

Grade:

1 Answers

Ravi
9 years ago
Start by writing Sin(179) as sin(180-1)

and sin(180-1)=sin1

Combine them to get (sin 1. sin 2..........sin 89)2 sin 90
sin 89=sin(90-1)=cos 1

Use this to obtain sin 2x and proceed similarily to combine the corresponding terms from both ends of the series to get the answer.